# MTTR-A: Measuring Cognitive Recovery Latency in Multi-Agent Systems

### v3

@@ Abstract
Ensuring cognitive stability in autonomous multi-agent systems (MAS) is a central challenge for large-scale, distributed AI. While existing observability tools monitor system outputs, they cannot quantify how rapidly agentic workflows recover once reasoning coherence has been lost. We adapt classical reliability metrics-Mean Time-to-Recovery (MTTR), Mean Time Between Failures (MTBF), and related ratios-into the cognitive domain, defining MTTR-A (Mean Time-to-Recovery for Agentic Systems) as a runtime measure of cognitive recovery latency. MTTR-A captures the time required for a MAS to detect reasoning drift and restore coherent operation, focusing on recovery of cognition rather than infrastructural repair.
A benchmark simulation using the AG News corpus and the LangGraph orchestration framework evaluates recovery latencies across reflex modes. Automated reflexes restore stability within approximately six seconds, while human-approval interventions require roughly twelve seconds. Across 200 runs, the median MTTR-A was $6.21\pm 2.14$ seconds, with MTBF $\approx 6.73\pm 2.14$ seconds and a normalized recovery ratio (NRR) of approximately 0.08.
By formalizing recovery latency as a measurable runtime property and deriving reliability bounds linking recovery time to cognitive uptime, this work establishes a foundation for runtime dependability in agentic AI systems. This work does not propose a new agent architecture or learning algorithm; instead, it focuses on system-level metricization and runtime dependability of existing agentic workflows.
@@ I Introduction
As multi-agent systems (MAS) built on large language models (LLMs) become increasingly autonomous and interconnected, such cognitive faults cannot be captured by conventional observability or uptime metrics [zhou2023autogen, amodei2016safety]. Recent surveys highlight similar reliability and coordination challenges in LLM-based MAS, underscoring the need for runtime governance and recovery mechanisms [han2025llmmaschallenges, guo2024llmmasurvey].
@@
To our knowledge, no prior orchestration framework defines a quantitative runtime metric for cognitive recovery; existing approaches focus on observability or schema enforcement, not on recovery latency itself [sun2024openagents, wang2023llmagentsurvey]. Industry reports concentrate on model capability and alignment rather than orchestration-level recovery, highlighting the absence of runtime reliability metrics in production AI systems [openai_gpt4, gemini_report, bai2022constitutional].
@@ I-A Our Approach: A Reliability Metricization Framework
This paper positions MTTR-A as a reliability metric for MAS. Rather than evaluating semantic accuracy or reasoning quality, we focus on runtime dependability -how quickly and predictably a distributed agent workflow detects, isolates, and recovers from cognitive faults. It unifies a reflex taxonomy, a latency-based metric family (MTTR-A, MTBF, NRR), and a LangGraph-based telemetry pipeline into a reproducible methodology for benchmarking agentic reliability.
As illustrated in Figure 2, this hierarchical reliability structure treats each agent as a subsystem composed of reflexive recovery components (rollback, replan, retry, approve), while the orchestration layer governs overall system-level dependability.
Multi-Agent System (MAS) System-level Cognitive ReliabilityAgent A (Subsystem)Agent B (Subsystem)Agent C (Subsystem)rollbackreplanretryapproveComponent-level reflexesrollbackreplanretryapproveComponent-level reflexesrollbackreplanretryapproveComponent-level reflexes
Figure 2: Hierarchical reliability model (illustrative). Each agent acts as a subsystem composed of reflexive recovery components (rollback, replan, retry, approve). For simplicity, three representative agents (A-C) are shown; the approach generalizes to any number of agents.
@@ II Related Work
LLM Agents and Deliberation: Chain-of-Thought [wei2022cot], ReAct [yao2022react], Toolformer [schick2023toolformer], and Tree-of-Thoughts [yao2023tot] provide prompting, acting, and tool-use patterns for single agents. Beyond single-agent prompting, multi-agent lines-Generative Agents [park2023generative], CAMEL [li2023camel], Reflexion [shinn2023reflexion], and Voyager [wang2023voyager]-together with recent surveys [wang2023llmagentsurvey, xi2023agentsurvey] map collaboration benefits and failure modes (e.g., coordination errors, unsafe tool calls, memory divergence). These works primarily improve ex-ante deliberation and coordination; they do not quantify ex-post recovery latency or runtime dependability once reasoning has drifted.
Recent studies have expanded MAS collaboration paradigms through reflection and debate mechanisms -such as Reflective Multi-Agent Collaboration [copper2024reflective], Chain-of-Agents [chainagents2024], MAGIS [magis2024], Multi-LLM Debate [multillmdebate2024], and DigiRL for device-control agents [digirl2024] - as well as empirical analyses on data-scale effects for UI-control agents [uicontrol2024]. While these frameworks enhance coordination quality and communication efficiency, they remain performance-oriented and do not formalize runtime dependability or recovery-latency metrics as addressed in the present work.
Resilient and Secure MAS: Decades of MAS work study consensus, fault tolerance, and security: resilient consensus under DoS and Byzantine settings [liu2023secureconsensus, wang2023resilientconsensus], fault-tolerant control [gao2022ftcc, liu2023ftcc, shi2024robust, sheng2024formation], adaptive/secure consensus [liu2023secureconsensus], and broader resilient modelling [varga2020resilient]. These works provide metrics, models, and control laws but do not standardize runtime recovery primitives for LLM-centric agent orchestration. Comprehensive reviews such as [guo2024llmmasurvey, han2025llmmaschallenges] have catalogued the rapid development of LLM-based multi-agent frameworks and identified persistent challenges in coordination, memory consistency, and agent orchestration. These analyses reinforce the motivation for a quantitative dependability metric - directly measures recovery dynamics rather than task performance or interaction quality.
@@
Recent orchestration frameworks such as AutoGen [zhou2023autogen], OpenAgents [sun2024openagents], and emerging control-plane studies including ALAS [alas_2025] and Advancing Agentic Systems [adv_agentic_2024] extend LLM coordination to multi-agent workflows, but they primarily address communication and execution graphs rather than quantitative recovery metrics such as MTTR-A.
Complementary benchmarking efforts such as Spider2-V [spider2v2024] evaluate multimodal professional-workflow agents, yet omit explicit reliability or recovery-time indicators, underscoring the absence of quantitative runtime dependability benchmarks in current agent evaluations.
@@ III-B Runtime Policy and Coordination Dynamics
The decision process that governs reflex selection is depicted in Figure 4, which maps fault triggers-such as tool errors, low confidence, or reasoning drift-to their corresponding recovery actions within the control loop.
@@
Trigger detected Tool error? Low confidence? Drift observed? tool-retrypasshuman-approvepassdrift-rollbackauto-replan
Figure 4: Policy decision tree used to select reflex actions. Branches correspond to fault types (tool error, low confidence, or drift) and map each to specific recovery primitives such as tool-retry or drift-rollback.
@@ IV-B Formal Reliability Metricization Framework
Reflex: tool-retryReflex: auto-replanReflex: rollbackAgent A Reasoning-Drift-Recovery $MTTR\text{-}A_{A},\;MTBF_{A}$Agent B Reasoning-Drift-Recovery $MTTR\text{-}A_{B},\;MTBF_{B}$Agent C Reasoning-Drift-Recovery $MTTR\text{-}A_{C},\;MTBF_{C}$Multi-Agent System (MAS) $MTTR\text{-}A_{sys},\;MTBF_{sys},\;NRR=1-\dfrac{MTTR\text{-}A_{sys}}{MTBF_{sys}}$
Figure 5: Hierarchical reliability model for a MAS. Agents A-C represent reasoning subsystems, each characterized by local recovery and stability metrics. System-level dependability emerges from the aggregation of agent-level performance.
@@
Figure 6: Decomposition of MTTR-A. Conceptual example showing the relative contributions of detection ($T_{\mathrm{detect}}$), decision ($T_{\mathrm{decide}}$), and execution ($T_{\mathrm{execute}}$) latencies as defined in Equation (5). This figure illustrates the theoretical structure of recovery latency.
@@ V Results and Discussion
Figure 10 visualizes the rolling 20-run MTTR-A average over the 200 iterations. The system exhibits steady recovery performance without trend escalation, suggesting stable reflex control and consistent orchestration efficiency throughout the experiment. Minor oscillations correspond to stochastic differences in recovery mode selection.
Figure 10: Rolling MTTR-A across 200 runs. The blue line represents a 20-run moving average; the dashed red line marks the overall median. No upward drift or performance degradation is observed.
V-A Latency Composition by Reflex Type
Figure 11 decomposes MTTR-A into three additive components: detection, decision, and execution latency. Execution time dominates recovery cost across all reflexes, especially in human-approval cases. The low decision latency confirms that policy selection overhead within LangGraph is negligible compared to the actual reflex execution duration.
Figure 11: Latency Composition by Reflex Type. Left: occurrence frequency of each recovery mode across all runs. Right: mean detection ($T_{\text{detect}}$), decision ($T_{\text{decide}}$), and execution ($T_{\text{execute}}$) times per mode. Execution dominates recovery cost, particularly in human-approval cases.
Automated reflexes (auto-replan and tool-retry) recover stability within 5-7 s, maintaining low variance and high throughput. Human oversight introduces expected governance latency but ensures compliance control.

### v4

@@ Abstract
Reliability in multi-agent systems (MAS) built on large language models is increasingly limited by cognitive failures rather than infrastructure faults. Existing observability tools describe failures but do not quantify how quickly distributed reasoning recovers once coherence is lost. We introduce MTTR-A (Mean Time-to-Recovery for Agentic Systems), a runtime reliability metric that measures cognitive recovery latency in MAS. MTTR-A adapts classical dependability theory to agentic orchestration, capturing the time required to detect reasoning drift and restore coherent operation. We further define complementary metrics, including MTBF and a normalized recovery ratio (NRR), and establish theoretical bounds linking recovery latency to long-run cognitive uptime. Using a LangGraph-based benchmark with simulated drift and reflex recovery, we empirically demonstrate measurable recovery behavior across multiple reflex strategies. This work establishes a quantitative foundation for runtime cognitive dependability in distributed agentic systems.
@@ I Introduction
As multi-agent systems (MAS) built on large language models (LLMs) become increasingly autonomous and interconnected, such cognitive faults cannot be captured by conventional observability or uptime metrics. Recent surveys highlight similar reliability and coordination challenges in LLM-based MAS, underscoring the need for runtime governance and recovery mechanisms [han2025llmmaschallenges, guo2024llmmasurvey].
@@
To our knowledge, no prior orchestration framework defines a quantitative runtime metric for cognitive recovery; existing approaches focus on observability or schema enforcement, not on recovery latency itself [sun2024openagents]. Industry reports concentrate on model capability and alignment rather than orchestration-level recovery, highlighting the absence of runtime reliability metrics in production AI systems [bai2022constitutional].
@@ I-A Our Approach: A Reliability Metricization Framework
This paper positions MTTR-A as a reliability metric for MAS. Rather than evaluating semantic accuracy or reasoning quality, we focus on runtime dependability -how quickly and predictably a distributed agent workflow detects, isolates, and recovers from cognitive faults. It unifies a reflex taxonomy, a latency-based metric family (MTTR-A, MTBF, NRR), and a LangGraph-based telemetry pipeline into a reproducible methodology for benchmarking agentic reliability. This hierarchical reliability structure treats each agent as a subsystem composed of reflexive recovery components (rollback, replan, retry, approve), while the orchestration layer governs overall system-level dependability.
@@ II Related Work
LLM Agents and Deliberation: Chain-of-Thought, ReAct, Toolformer, and Tree-of-Thoughts [yao2023tot] provide prompting, acting, and tool-use patterns for single agents. Beyond single-agent prompting, multi-agent lines-Generative Agents, CAMEL, Reflexion [shinn2023reflexion], and Voyager [wang2023voyager]-together with recent surveys [xi2023agentsurvey] map collaboration benefits and failure modes (e.g., coordination errors, unsafe tool calls, memory divergence). These works primarily improve ex-ante deliberation and coordination; they do not quantify ex-post recovery latency or runtime dependability once reasoning has drifted.
Recent studies have expanded MAS collaboration paradigms through reflection and debate mechanisms -such as Reflective Multi-Agent Collaboration [copper2024reflective], Chain-of-Agents [chainagents2024], and MAGIS [magis2024].
While these frameworks enhance coordination quality and communication efficiency, they remain performance-oriented and do not formalize runtime dependability or recovery-latency metrics as addressed in the present work.
Resilient and Secure MAS: Decades of MAS work study consensus, fault tolerance, and security: resilient consensus under DoS and Byzantine settings [liu2023secureconsensus, wang2023resilientconsensus], fault-tolerant control [gao2022ftcc], adaptive/secure consensus [liu2023secureconsensus], and broader resilient modelling [varga2020resilient]. These works provide metrics, models, and control laws but do not standardize runtime recovery primitives for LLM-centric agent orchestration.
These analyses reinforce the motivation for a quantitative dependability metric - directly measures recovery dynamics rather than task performance or interaction quality.
@@
Recent orchestration frameworks such as AutoGen, OpenAgents, and emerging control-plane studies including ALAS and Advancing Agentic Systems [adv_agentic_2024] extend LLM coordination to multi-agent workflows, but they primarily address communication and execution graphs rather than quantitative recovery metrics such as MTTR-A.
Complementary benchmarking efforts evaluate multimodal professional workflow agents, yet omit explicit reliability or recovery-time indicators, underscoring the absence of quantitative runtime dependability benchmarks in current agent evaluations.
@@ III-B Runtime Policy and Coordination Dynamics
The decision process that governs reflex selection maps fault triggers-such as tool errors, low confidence, or reasoning drift-to their corresponding recovery actions within the control loop.
@@ V Results and Discussion
To assess temporal stability, we analyzed the rolling 20-run MTTR-A average over the 200 experimental iterations. The results indicate steady recovery performance without systematic escalation, suggesting stable reflex control and consistent orchestration efficiency throughout the experiment. Minor fluctuations are attributable to stochastic variation in recovery mode selection rather than degradation in system behavior.
To analyze recovery behavior, MTTR-A was decomposed into three additive components: detection, decision, and execution latency. Across all reflex modes, execution time constitutes the dominant contribution to recovery cost, with the effect most pronounced in human-approval cases. Decision latency remains consistently low, indicating that policy selection overhead within the LangGraph orchestration layer is negligible relative to reflex execution time.
Fully automated reflexes, such as auto-replan and tool-retry, restore stability within approximately 5–7 s with low variance, supporting high-throughput operation. Human-in-the-loop interventions introduce expected governance latency but provide necessary compliance and control in high-risk scenarios.
@@ Appendix C Experimental Configuration
[Uncaptioned image] Barak Or (Member, IEEE) received a B.Sc. degree in aerospace engineering (2016), a B.A. degree (cum laude) in economics and management (2016), and an M.Sc. degree in aerospace engineering (2018) from the Technion–Israel Institute of Technology. He graduated with a Ph.D. degree from the University of Haifa, Haifa (2022). His research interests include navigation, deep learning, sensor fusion, and estimation theory.